The area of a rug is 108 square feet and the length of its diagonal is 15 feet. Find the length and width of the rug.
Answer by scott8148(6628)

let x=length and y=width ___ xy=108 ___ x^2+y^2=15^2

y=108/x ___ x^2+108^2/x^2=15^2

x^4+108^2=225x^2 ___ x^4-225x^2+11664=0 ___ (x^2-144)(x^-81)=0 ___ (x+12)(x-12)(x+9)(x-9)=0